OpenStudy (anonymous):

The hypotenuse of a 30- 60- 90 triangle has length 15. What is the length of the side opposite the 60 angle? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

sin 60 = sqrt3 /2 = opp side / 15 2 * opp side = 15*sqrt3 opp side = 15*sqrt3 / 2

OpenStudy (anonymous):

I haven't got a calculator handy to ]finish this off!!

OpenStudy (anonymous):

12.99 thanks

OpenStudy (anonymous):

12.99

OpenStudy (anonymous):

\[15\text{Sin}[\pi /3]=\frac{15 \sqrt{3}}{2}=12.9904 \]